OpenStudy (anonymous):

How would I solve this linear inequality and graph it? 3x-y<4

OpenStudy (e.mccormick):

Solve it for y. Then it is pretty similar to graphing a line. But you also end up with a shaded area.

OpenStudy (anonymous):

You'd put it as y=-3x+4, correct? That's about as far as I've gotten

OpenStudy (e.mccormick):

ummm.... how did your, ah. You dropped the - on the y and forgot about it.

OpenStudy (anonymous):

That's why I'm asking for help. I don't really know what I'm doing

OpenStudy (e.mccormick):

Well, what you really have at this point is: \(-y<-3x+4\) If you divide out the -, or -1, what happens to that?

OpenStudy (anonymous):

y<3x-4?

OpenStudy (e.mccormick):

Very, very close! When you divide/multiply a negative in an inequality, the arrow changes! \(< \rightarrow >\)

OpenStudy (e.mccormick):

So: \(y>3x-4\) c: Do you get that part?

OpenStudy (anonymous):

yup so what would be the next step?

OpenStudy (e.mccormick):

As an example: \(5>-1\) \((-1)(5>-1)\) \(-5<1\) OK, so next, you need to find that line. Now,there is a clue as to if a line is part of the graph, or if it is dashed.... or are you doing these on number lines? On a number line it is an open circle or a closed one.

OpenStudy (anonymous):

I think that < is just a dashed line but if it was greater/less than or equal to it would be solid. Or is it the other way around?

OpenStudy (e.mccormick):

You got it! \(=\), \(\le\), and \(\ge\) are solid lines and part of the graphed value. \(\ne\), \(<\), and \(>\) are dashed lines and the border next to the graphed value.

OpenStudy (e.mccormick):

OK, now, you goow with the graphing the line part?

OpenStudy (e.mccormick):

good...

OpenStudy (anonymous):

I don't know what I'm supposed to plug in, in order to get the points

OpenStudy (e.mccormick):

0 is easy. Put 0 in for y and get the x intercept. Then put in 0 for x and get the y intercept.

OpenStudy (anonymous):

and those are the only points I'd need?

OpenStudy (e.mccormick):

Any two points determine a line. Graph them and you can draw a dashed line through them. Line part done and all that is left is shading!

OpenStudy (e.mccormick):

For the shading: \(y \le stuff\) and \(y < stuff\) mean the y values are valid below the line. \(y\ge stuff\) and \(y> stuff\) mean the y values are valid above the line. On a side note, why I picked 0: The two most important numbers in doing math problems: 0 and 1. (\(e\text{ and } \pi\) are great, but 0 and 1 rule!) See, \(x+0 = x+a-a\) which comes in handy for completing the square. \(\dfrac{x}{n}=\dfrac{x}{n}\times \dfrac{a}{a}=\dfrac{xa}{an}\) which is just multiply by 1 and comes in handy for tons of things!

OpenStudy (e.mccormick):

Did your points come out something like this? 

OpenStudy (anonymous):

Yeah they came out that way and I shaded above where the line would be and got the problem right :) thank you so much for taking the time to explain things to me

OpenStudy (e.mccormick):

So this: https://www.desmos.com/calculator/2vme81vjwq Good! have fun!